# General surgery vision transformer: A video pre-trained foundation model for general surgery

### v2

@@ 2.1 Foundation models in medicine
In medicine, foundations models have been developed for applications such as radiology, endoscopic video analysis, and x-ray diagnoses. These foundation models can largely be categorized into three architectures: language models, vision models, and vision-language models. Language models are used in settings where diagnoses occur through dialogue, such as case studies presented from medical exams [Sin+23, Sch+24, ZS23, Nor+23] or extracting data from clinical notes [Agr+22]. Vision models are useful for medical tasks requiring visual perception, such as medical segmentation [Che+21a, Ma+24] (e.g. organs, tumors, medical tools) or disease classification [Zho+23]. Vision-language models combine both vision and language, enabling the interaction of visual input with natural language, such as asking questions about images of a disease [Hua+23, Li+23]. Vision-based foundation models could be particularly impactful, as medical images are high dimensional and more challenging to acquire at volume, with the foundation model enabling medical tasks to be solved with less data.
@@ 2.3 Surgical Datasets
Tool classification has some of the largest datasets in surgery [Twi+16, Al ̵+19], ranging from 86k to 900k frames. There also exists a variety of surgical datasets focused on tool segmentation from various operations [Szn+12, All+20, Hon+20], typically ranging in size from 100-10K images. Perhaps the largest surgical dataset is SurgToolLoc with 44M frames [Zia+23], focusing on tool presence classification. Our work focuses on unsupervised pre-training, which most closely related to work done in endoscopy [Bat+23] and with MRI data [Pra+24]. We note that these models were trained on relatively small datasets compared to our work (307 and 700K images compared to 70M images).
@@ 3.1 GSViT: General Surgery Vision Transformer
The architecture we chose for GSViT is based on the EfficientNet ViT structure [Liu+23]. The design of this model emphasizes speed in inference, minimizing memory and computation during inference and training time. We believe this is important for surgical applications, where algorithms run in real-time. Below, we describe this architecture in detail (also see Fig. 1):
@@ 3.1.2 Pre-training through video prediction
Unlike the applications of other medical foundation models which were trained for image reconstruction, surgery is spatial and temporal. The majority of image-pretraining techniques, such as direct image reconstruction or masking techniques, train on a single image at an instance in time. However, we believe that the latent space embedding for surgical problems should capture both spatial and temporal properties and require developing models that account for both space and time. Recent work has explored pre-training using video prediction [Gup+22], showing improved results for spatio-temporal problems. We take a similar approach GSViT, performing next-frame reconstruction for pre-training with a 1 frame input and a 1 second prediction (1 frame). We found that the addition of masking, such as in [Gup+22], did not improve performance and opted for the next-frame reconstruct as demonstrated in [Gao+22].
@@ 4.1 Creating GenSurgery from public YouTube videos
Further, each video was categorized into the specific operation that was performed as well as whether it was a laparoscopic or robotic surgery. The number of videos for each type of operation is demonstrated visually in Figure 2, as well the number of frames total for each type of operation. We note that the operation type with the largest number of frames versus total videos differs quite significantly because operations can vary significantly in time (e.g. hernia repair averages 90.8 minutes [Qab+21] and hepatectomy averages 424.4 minutes [JKK22]). In total, GenSurgery has a total of 70M frames.
@@ 4.3 GSViT can run in real-time applications
We compare the runtime inference time (no gradient accumulation) for each the of following models on input images from the Cholec80 dataset. Here, we note that GSViT is able to process 10621 images per second [Liu+23] which is 10.6 images per millisecond. In a non-parallelized process we attain a runtime average of $12.1\pm 0.1$ ms on a single 12GB NVIDIA RTX A5500 GPU. We can compare this with two state-of-the-art efficient architectures, EfficientNet-B0 [TL19] and GLiT-Tiny [Che+21], which attain a throughput average of $4532$ and $3516$ images per second (inference time of $4.5$ and $3.5$ images per millisecond) respectively.
@@ 4.4 Surgical phase detection
We use a learning rate of $43e-$ 4 with exponential decay and a batch size of $128$. The MLP classifier head outputs a $7$ dimensional classification vector and has two layers with $2048$ and $512$ hidden neurons respectively. The hidden neurons have exponential linear unit (ELU) activations [CUH15] followed by a layer normalization [BKH16].
@@
The results of GSViT-Cholecystectomy trained on Cholec80 is shown in Table 1. We demonstrate comparable performance to previous models, obtaining $86.3\%$ accuracy, at a significantly reduced computational cost and a reduced amount of tuneable parameters [^1] The total number of parameters is 26.1M, with 13.7M tuneable parameters and an additional 12.4M non-tuneable parameters for the foundation model. We also note that our approach recognizes phase by only looking at the current frame, whereas other approaches such as TeCNO [Cze+20] take in the entire video history for a phase classification. Considering this, our model is the highest performance single-frame model. We incorporate this to make it more feasible for real-time surgical applications.
@@ 5 Discussion
Future work could expand the pre-training by incorporating temporal and spatial mask reconstruction. Additionally, methods for obtaining higher performance on the Cholec80 dataset could be explored such as temporal convolutions [Cze+20], long videos [Liu+23a] or recurrent convolutions [Jin+20]. We also note that our dataset could be used for building a more general-purpose medical foundation model that includes surgery as well as additional modalities such as MRI and CT images [Sch+24a].

### v3

@@ 2.1 Foundation models in medicine
In medicine, foundations models have been developed for applications such as radiology, endoscopic video analysis, and x-ray diagnoses. These foundation models can largely be categorized into three architectures: language models, vision models, and vision-language models. Language models are used in settings where diagnoses occur through dialogue, such as case studies presented from medical exams [singhal2023towards, schmidgall2024addressing, ziaei2023language, nori2023can] or extracting data from clinical notes [agrawal2022large]. Vision models are useful for medical tasks requiring visual perception, such as medical segmentation [chen2021transunet, ma2024segment] (e.g. organs, tumors, medical tools) or disease classification [zhou2023foundation]. Vision-language models combine both vision and language, enabling the interaction of visual input with natural language, such as asking questions about images of a disease [huang2023visual, li2023llava]. Vision-based foundation models could be particularly impactful, as medical images are high dimensional and more challenging to acquire at volume, with the foundation model enabling medical tasks to be solved with less data.
@@ 2.3 Surgical Datasets
Tool classification has some of the largest datasets in surgery [twinanda2016endonet, al2019cataracts], ranging from 86k to 900k frames. There also exists a variety of surgical datasets focused on tool segmentation from various operations [sznitman2012data, allan20202018, hong2020cholecseg8k], typically ranging in size from 100-10K images. Perhaps the largest surgical dataset is SurgToolLoc with 44M frames [zia2023surgical], focusing on tool presence classification. Our work focuses on unsupervised pre-training, which most closely related to work done in endoscopy [batic2023whether] and with MRI data [prabhakar2024vit]. We note that these models were trained on relatively small datasets compared to our work (307 and 700K images compared to 70M images).
@@ 3.1 GSViT: General Surgery Vision Transformer
The architecture we chose for GSViT is based on the EfficientNet ViT structure [liu2023efficientvit]. The design of this model emphasizes speed in inference, minimizing memory and computation during inference and training time. We believe this is important for surgical applications, where algorithms run in real-time. Below, we describe this architecture in detail (also see Fig. 1):
@@ 3.1.2 Pre-training through video prediction
Unlike the applications of other medical foundation models which were trained for image reconstruction, surgery is spatial and temporal. The majority of image-pretraining techniques, such as direct image reconstruction or masking techniques, train on a single image at an instance in time. However, we believe that the latent space embedding for surgical problems should capture both spatial and temporal properties and require developing models that account for both space and time. Recent work has explored pre-training using video prediction [gupta2022maskvit], showing improved results for spatio-temporal problems. We take a similar approach GSViT, performing next-frame reconstruction for pre-training with a 1 frame input and a 1 second prediction (1 frame). We found that the addition of masking, such as in [gupta2022maskvit], did not improve performance and opted for the next-frame reconstruct as demonstrated in [gao2022simvp].
@@ 4.1 Creating GenSurgery from public YouTube videos
Further, each video was categorized into the specific operation that was performed as well as whether it was a laparoscopic or robotic surgery. The number of videos for each type of operation is demonstrated visually in Figure 2, as well the number of frames total for each type of operation. We note that the operation type with the largest number of frames versus total videos differs quite significantly because operations can vary significantly in time (e.g. hernia repair averages 90.8 minutes [qabbani2021robotic] and hepatectomy averages 424.4 minutes [jang2022early]). In total, GenSurgery has a total of 70M frames.
@@ 4.3 GSViT can run in real-time applications
We compare the runtime inference time (no gradient accumulation) for each the of following models on input images from the Cholec80 dataset. Here, we note that GSViT is able to process 10621 images per second [liu2023efficientvit] which is 10.6 images per millisecond. In a non-parallelized process we attain a runtime average of $12.1\pm 0.1$ ms on a single 12GB NVIDIA RTX A5500 GPU. We can compare this with two state-of-the-art efficient architectures, EfficientNet-B0 [tan2019efficientnet] and GLiT-Tiny [chen2021glit], which attain a throughput average of $4532$ and $3516$ images per second (inference time of $4.5$ and $3.5$ images per millisecond) respectively.
@@ 4.4 Surgical phase detection
We use a learning rate of $43e-$ 4 with exponential decay and a batch size of $128$. The MLP classifier head outputs a $7$ dimensional classification vector and has two layers with $2048$ and $512$ hidden neurons respectively. The hidden neurons have exponential linear unit (ELU) activations [clevert2015fast] followed by a layer normalization [ba2016layer].
@@
The results of GSViT-Cholecystectomy trained on Cholec80 is shown in Table 1. We demonstrate comparable performance to previous models, obtaining $86.3\%$ accuracy, at a significantly reduced computational cost and a reduced amount of tuneable parameters [^1] The total number of parameters is 26.1M, with 13.7M tuneable parameters and an additional 12.4M non-tuneable parameters for the foundation model. We also note that our approach recognizes phase by only looking at the current frame, whereas other approaches such as TeCNO [czempiel2020tecno] take in the entire video history for a phase classification. Considering this, our model is the highest performance single-frame model. We incorporate this to make it more feasible for real-time surgical applications.
@@ 5 Discussion
Future work could expand the pre-training by incorporating temporal and spatial mask reconstruction. Additionally, methods for obtaining higher performance on the Cholec80 dataset could be explored such as temporal convolutions [czempiel2020tecno], long videos [liu2023lovit] or recurrent convolutions [jin2020multi]. We also note that our dataset could be used for building a more general-purpose medical foundation model that includes surgery as well as additional modalities such as MRI and CT images [schmidgall2024general, schmidgall2024robots].